19 August 2015: Wednesday

LDAP:

  • LDAP Update:
    • One issue of a new employee showing up to replace an existing employee, but the order was add,delete instead of delete,add so the update code threw a fit.  Manually adding the entry (the delete happened just fine after the fit).
  • ACCOUNT Update:
    • No issues.
  • ACLs:
    • Another batch of ACL changes for openstack.

Backups:

  • Save Groups:
    • Failed:
      • cmsi-noah decided not to talk to us last night.
      • windowpane (as usual)
  • Clones:
    • Tape clones continue to try to complete.
  • Vault Tape Reclamation:
    • Got one (1) back.
  • Problems:
    • Need to swap the location of the two HBA’s in stornode5 to see if the problem follows the card.  I will have to come back later to do that as the backups going through it have just not completed yet.
      • Came back at 5:30 – swapped the cards.
        • card lights being on solid (not flashing) followed the card
        • However, the card that works in riser 1 did not work in riser 2 (the lights were right, but no data was passed).  May have a fried card and riser or mother board socket /sigh

VPN Change:

  • Testing the one-shot update code.  Bugs!  I need a new can of RAID.
  • Gotem!  Another round of testing to be certain.
  • It is live.

Tags: ,

Leave a Reply

You must be logged in to post a comment.